log(6)x+log(6)(x+16)=2
I dont know how to solve this, and im supposed to check for extraneous solutions

Answer by scott8148(6628)

adding logs means multiplying their arguments

x(x + 16) = 6^2 ___ x^2 + 16x = 36 ___ x^2 + 16x - 36 = 0

factoring ___ (x + 18)(x - 2) = 0